# thanks: Corresponding author

### v3

@@ VI Discussion and Conclusions
We have studied the spin polarization distribution and evolution in dark matter halos in fully non-linear, nonrelativistic, numerical simulations of vector dark matter (VDM). Our work builds on recent studies such as Refs. Chen et al. (2023); Jain et al. (2023a); Amaral et al. (2024). We simulated different stochastic initial conditions, and studied the spin polarization in the regime of time scales much longer than the coherence time. We found that non-linear gravitational evolution does not change the spin distribution in the VDM halo on average compared to the initial conditions, due to conservation laws being respected. The spin in local regions, however, undergoes fluctuations, with a stronger amplitude of fluctuation in the outer regions of the halo compared to near the core. The core has large spin in all initial conditions.
@@
It is possible to have still more general polarization settings using the correlation matrix formalism we presented in Ref. Chen et al. (2023). Such a setting can tune the desired value of $S^{i}$ in any component, and thus the value of $\kappa$ relevant for experiment (the case with very small $S/N$ is generated spontaneously by numerical effects in our CF simulations). Such a random mixed state model is generated spontaneously by our random initial conditions from the power spectrum.
@@ Acknowledgements
We especially thank Xiaolong Du for providing significant assistance during our research. We thank Lei Pan, Chao Gao, Yikun Gu, Hongyi Zhang, ZhiPan Li and Mudit Jain for beneficial discussions.

### v4

@@ VI Discussion and Conclusions
We have studied the spin polarization distribution and evolution in dark matter halos in fully non-linear, nonrelativistic, numerical simulations of vector dark matter (VDM). Our work builds on recent studies such as Refs. Chen et al. (2023a); Jain et al. (2023a); Amaral et al. (2024). We simulated different stochastic initial conditions, and studied the spin polarization in the regime of time scales much longer than the coherence time. We found that non-linear gravitational evolution does not change the spin distribution in the VDM halo on average compared to the initial conditions, due to conservation laws being respected. The spin in local regions, however, undergoes fluctuations, with a stronger amplitude of fluctuation in the outer regions of the halo compared to near the core. The core has large spin in all initial conditions.
@@
It is possible to have still more general polarization settings using the correlation matrix formalism we presented in Ref. Chen et al. (2023a). Such a setting can tune the desired value of $S^{i}$ in any component, and thus the value of $\kappa$ relevant for experiment (the case with very small $S/N$ is generated spontaneously by numerical effects in our CF simulations). Such a random mixed state model is generated spontaneously by our random initial conditions from the power spectrum.
@@ Acknowledgements
We thank Xiaolong Du for providing significant assistance during our research, and collaboration in an early stage of this work. We thank Lei Pan, Chao Gao, Yikun Gu, Hongyi Zhang, ZhiPan Li and Mudit Jain for beneficial discussions, and Mudit Jain for reading an early version of this manuscript.